Pamela has a garden 10 m by 17 m. She uses part of it for a badminton court which is inthe shape of a rectangle by subtracting equal amounts from both the length and the width. The area of the garden is now 92 sq. meters. Find the dimensions of the badminton court. Please Help. I have no idea how to set this up.
Answer by mananth(16946)

Total area of garden =17*10 = 170 m^2
Equal amounts in length are subtracted from the sides . -----x
Width of court = 10-2x
length of court = 17-2x
Area of court = (10-2x)(17-2x)
Area of garden - area of court = 92
170-(10-2x)(17-2x)= 92
170-(170-20x-34x+4x^2)=92
170-170+54x-4x^2=92
OR
4x^2-54x+92=0
4x^2-8x-46x+82=0
4x(x-2)-46(x-2)=0
(4x-46)(x-2)=0
x=2 OR 46/4
46/4 is not possible
width = 2 m
Dimensions are 10-4 =6m
17-4 = 13m